# On Inverse Problems, Parameter Estimation, and Domain Generalization

### v1

@@ 4 Discrete parameter estimation
    $\theta-X-Y-\hat{X}-\hat{\theta}_{\hat{X}}$ (13)
where $\hat{X}=\mathcal{F}(Y)$, where the mapping $\mathcal{F}:\mathcal{Y}^{d\times 1}\rightarrow\mathcal{X}^{n\times 1}$ is an estimator of $X$ (reconstruction) given the observation $Y$ (or $m$ examples), $\hat{\theta}_{Y}=g(Y)$ and $\hat{\theta}_{\hat{X}}=f(\hat{X})$ are predictions of $\theta$. Note that in some cases the estimation $\hat{X}$ may be obtained from an example (or examples) of $Y|\theta$, that is of a measurement(s) belonging to a specific $\theta$, without the knowledge of $\theta$, but with the knowledge that all these examples originate in the same $\theta$. Alternatively, one could recover $\hat{X}$, from $Y$, based on a recovery method that is general to all $Y$’s regardless of the $\theta$ they originate from. The question at the core of this research paper is: Which path leads to a better $\theta$ estimation and when? Or, in other words, what would be the optimal path towards parameter estimation related with inverse problems?
We note that $Y$ is not necessarily a deterministic function of $X$. For example, in an inverse problem, $Y$ could be a degraded signal originating in $X$, where the relationship between $X$ and $Y$ could be linear or non-linear, with or without additive noise, such that generally $y=u(x)+v(x)$, where $u(\cdot)$ and $v(\cdot)$ depend on $x$, and are deterministic or random functions. For example, $u(\cdot)$ could represent a blurring process, while $v(\cdot)$ represents an additive noise that could depend on the true value of $x$. Additionally, we consider two types of estimators for $\hat{X}$: deterministic and stochastic. Given a particular input $Y=y$, a restoration algorithm $\mathcal{F}:\mathcal{Y}^{d\times 1}\rightarrow\mathcal{X}^{n\times 1}$ produces an estimate $\hat{X}$ according to some distribution $p_{\hat{X}|Y}(\cdot|y)$ such that the estimate $\hat{X}$ is statistically independent of $X$ given $Y$ For deterministic algorithms, $p_{\hat{X}|Y}(\cdot|y)$ is a delta function for every $y$, while for stochastic algorithms it is a non-degenerate distribution. Furthermore, in many practical cases the estimated distribution $p_{\hat{X}|Y}(\cdot|y)$ is designed as a mapping that is independent of $\theta$.
@@ 4.1 Perfect Perception Estimator
Definition 4 (Perfect Perception Recovery)
@@
It has been established in previous work (Blau and Michaeli, 2018) that a perfect perceptual quality estimator obeys $p_{\hat{X}}(x)=p_{X}(x)$. In (Blau and Michaeli, 2018), Blau et al. (2018) also showed that perceptual quality and distortion are at odds with each other, and that for the square error distortion $\|\hat{x}-x\|^{2}$, perfect perceptual quality need not sacrifice more than 3dB in PSNR.
@@ 4.2 Theoretical Analysis
Consider M classes $\theta_{i}$ with the conditional probability distribution functions (pdf’s) $p(\mathbf{x}|\theta_{i})$ defined on the observation space $\mathcal{X}^{n}$, with priors $P(\theta_{i})$, respectively, $i=1,...,M$. The posterior pdf of $\theta_{i}$ is
$q_{i}(\mathbf{x})\triangleq p(\theta_{i}|\mathbf{x})=\frac{P(\theta_{i})p(\mathbf{x}|\theta_{i})}{p(\mathbf{x})},$ (14)
@@
According to Bayes rule each point $\mathbf{x}$ is assigned to the class $\hat{k}$ with the maximum posterior probability, $\hat{k}=\mathrm{argmax}_{k}\ q_{k}(x),\ k=1,...,M$ , or equivalently, the minimal loss function (Lissack and Fu, 1976). Namely, $\mathbf{x}$ is assigned to class $\theta_{i}$, if $\mathbf{x}\in\Gamma_{i}$, where
$\Gamma_{i}=\Big{\{}\mathbf{x}\in\mathcal{X}^{n}\ \Big{|}\ \sum^{M}_{k=1}c_{ik}P(\theta_{k})p(\mathbf{x}|\theta_{k})<\sum^{M}_{k=1}c_{lk}P(\theta_{k})p(\mathbf{x}|\theta_{k}),\ l=1,...,M,l\neq i\Big{\}},$ (15)
@@
$P(\theta_{1})p(\mathbf{x}|\theta_{1})<P(\theta_{2})p(\mathbf{x}|\theta_{2})\quad=>\quad\mathbf{x}\sim p_{\theta_{2}}(\mathbf{x}),$ (17)
@@
[…]
@@
$\Gamma_{1}=\Big{\{}\mathbf{x}\in\mathcal{X}^{n}\ \Big{|}\ P(\theta_{2})p(\mathbf{x}|\theta_{2})<P(\theta_{1})p(\mathbf{x}|\theta_{1})\Big{\}},$ (19)
@@
$\Gamma_{2}=\Big{\{}\mathbf{x}\in\mathcal{X}^{n}\ \Big{|}\ P(\theta_{1})p(\mathbf{x}|\theta_{1})<P(\theta_{2})p(\mathbf{x}|\theta_{2})\Big{\}}.$ (20)
Lissack and Fu (1976) defined a separability measure between two classes,
$J^{x}_{\alpha}(\theta_{1},\theta_{2})=E\phi_{\alpha}(\mathbf{x}),$ (21)
@@
$\phi_{\alpha}(\mathbf{x})=|q_{1}(\mathbf{x})-q_{2}(\mathbf{x})|^{\alpha}\quad 0\leq\alpha\leq\infty.$ (22)
@@
$P_{e}=\frac{1}{2}[1-J_{1}(\theta_{1},\theta_{2})].$ (23)
@@
[…]
The bounds for $\alpha\neq 1$ and for $M>2$ is provided in Lissack and Fu (1976). In other words, the probability of misrecognition is inherently bounded by the absolute value of the difference between the posterior pdfs.
@@
$J^{x}_{1}(\theta_{1},\theta_{2})=\int\big{|}P(\theta_{2})p(\mathbf{x}|\theta_{2})-P(\theta_{1})p(\mathbf{x}|\theta_{1})\big{|}d\mathbf{x}.$ (24)
@@
$J^{x}_{1}(\theta_{1},\theta_{2})=\int\big{|}\beta p(\mathbf{x}|\theta_{1})-(1-\beta)p(\mathbf{x}|\theta_{2})\big{|}d\mathbf{x}.$ (25)
@@
$J^{y}_{1}(\theta_{1},\theta_{2})=\int\big{|}\beta p(\mathbf{y}|\theta_{1})-(1-\beta)p(\mathbf{y}|\theta_{2})\big{|}d\mathbf{y}.$ (26)
@@
[…]
@@
$J^{y}_{1}(\theta_{1},\theta_{2})$ $=\int\Big{|}\beta\int p(\mathbf{y}|\mathbf{x})p(\mathbf{x}|\theta_{1})d\mathbf{x}-(1-\beta)\int p(\mathbf{y}|\mathbf{x})p(\mathbf{x}|\theta_{2})d\mathbf{x}\Big{|}d\mathbf{y}$ (28)
@@
[…]
Theorem 5 (Bayes error bound following signal transformation)
@@
$J^{\hat{\mathbf{x}}}_{1}(\theta_{1},\theta_{2})=\int\Big{|}\int p(\hat{\mathbf{x}}|\mathbf{y})\big{[}\beta p(\mathbf{y}|\theta_{1})d\mathbf{y}-(1-\beta)p(\mathbf{y}|\theta_{2})\big{]}d\mathbf{y}\Big{|}d\hat{\mathbf{x}}\\ \leq J^{y}_{1}(\theta_{1},\theta_{2})\leq J^{x}_{1}(\theta_{1},\theta_{2}).$ (30)
@@
Theorem 6 (Bayes error bound following class-agnostic signal recovery)
@@
$P_{e}(\hat{\mathbf{x}})\geq P_{e}(\mathbf{y})\geq P_{e}(\mathbf{x}).$ (31)
@@
Theorem 7 (Bayes error bound following perfect conditional perception recovery)
@@
$J^{\hat{x}}_{1}(\theta_{1},\theta_{2})=\int\Big{|}\beta\int p(\mathbf{x}|\theta_{1})p(\hat{\mathbf{x}}|\mathbf{x},\theta_{1})d\mathbf{x}-(1-\beta)\int p(\mathbf{x}|\theta_{2})p(\hat{\mathbf{x}}|\mathbf{x},\theta_{2})d\mathbf{x}\Big{|}d\hat{\mathbf{x}}.$ (33)
@@
$\arg\max_{\mathbf{x}}p_{\theta}(\mathbf{x},\mathbf{y})=\arg\max_{\mathbf{x}}p(\mathbf{x}|\mathbf{y})p(\theta|\mathbf{x})$ (34)
@@
Related Work. Jalal et al. (2021) state that in light of the “White Obama” controversy (Menon et al., 2020b), it has been suggested that reconstruction algorithms are biased because the datasets used for training and inference are unbalanced, thus the obtained models misrepresent the true population distribution. Jalal et al. (2021) also claim that reconstruction algorithms that are designed to maximize accuracy are inclined to increase bias. They define proportional representation (PR) such that given a collection of sets of images $\{\mathcal{X}_{\theta_{i}}\}_{i=1}^{K}$ where $\theta_{i}$ is a vector of sensitive attributes and each $\mathcal{X}_{\theta_{i}}$ represents the group carrying the sensitive attributes $\theta_{i}$,
$P(\hat{\mathbf{x}}\in\mathcal{X}_{\theta_{i}})=P(\mathbf{x}\in\mathcal{X}_{\theta_{i}})\ \forall i.$ (35)
@@
$P(\hat{\mathbf{x}}\in\mathcal{X}_{\theta_{i}}|\mathbf{y})=P(\mathbf{x}\in\mathcal{X}_{\theta_{i}}|\mathbf{y})\ \forall i.$ (36)
They show that randomly sampling from the posterior $\hat{\mathbf{x}}\sim p(\mathbf{x}|\mathbf{y})$ achieves CPR and PR. Hence, fairness is accomplished because every image that could have led to the measurement could be represented, and not just the most likely one. Note that by these definitions the reconstruction is stochastic, and not deterministic, and that any reconstruction that falls outside the set $\mathcal{X}_{\theta_{i}}$ is considered a wrong outcome. Furthermore, these definitions ignore the distribution $p_{\theta_{i}}(\mathbf{x})$. Thus, Ohayon et al. (Ohayon et al., 2024a) propose other measures of fairness to improve the estimation of the distribution of the different groups. However, note that, in our case study, our objective is to estimate $\theta_{i}$ to the best accuracy. We are not seeking to improve fairness nor to provide a more diverse distribution of the reconstruction $\mathbf{x}$.
@@ 4.3.1 Naive tree example
Consider the setting partially depicted in figure 7 such that, $\theta=\{\theta_{1},\theta_{2}\}$ represent two classes with probabilities $P(\theta_{1})=0.8,\ P(\theta_{2})=0.2$. The $x$ values in each class are $A_{X}=\{0,1,2\}$, $B_{X}=\{3,4,5\}$, which are conditionally equiprobable, $\big{\{}x\in A_{X}:p_{\theta_{1}}(x)=\big{[}\frac{1}{3},\frac{1}{3},\frac{1}{3}\big{]}\big{\}}$, $\big{\{}x\in B_{X}:p_{\theta_{2}}(x)=\big{[}\frac{1}{3},\frac{1}{3},\frac{1}{3}\big{]}\big{\}}$. The $y$ values in each class are $A_{Y}=\{0.5,1.5,2.5\}$, $B_{Y}=\{1.5,3.5,4.5\}$, and the conditional probabilities of observing $y=1.5$ are $p(y=1.5|x=4)=1,\ p(y=1.5|x=1)=0.4$. Namely, the original sets are disjoint but the observation sets partially coincide, and one set is 4 times more likely than the other.

### v2

@@ 4 Discrete parameter estimation
    (17) $\theta-X-Y-\hat{X}-\hat{\theta}_{\hat{X}}$
[…]
We note that $Y$ is not necessarily a deterministic function of $X$. For example, in an inverse problem, $Y$ could be a degraded signal originating in $X$, where the relationship between $X$ and $Y$ could be linear or non-linear, with or without additive noise, such that generally $y=u(x)+v(x)$, where $u(\cdot)$ and $v(\cdot)$ depend on $x$, and are deterministic or random functions. For example, $u(\cdot)$ could represent a blurring process, while $v(\cdot)$ represents an additive noise that could depend on the true value of $x$. Additionally, we consider two types of estimators for $\hat{X}$: deterministic and stochastic. In a continuous setting, given a particular input $Y=y$, a restoration algorithm $\mathcal{F}:\mathcal{Y}^{d\times 1}\rightarrow\mathcal{X}^{n\times 1}$ produces an estimate $\hat{X}$ according to some conditional probability density function $p_{\hat{X}|Y}(\cdot|y)$ such that the estimate $\hat{X}$ is statistically independent of $X$ given $Y$ . For deterministic algorithms, the conditional probability distribution $P_{X|Y}(\cdot|y)$ is a delta measure for almost every $y$, while for stochastic algorithms it is a non-degenerate distribution. Furthermore, in many practical cases the estimated distribution $p_{\hat{X}|Y}(\cdot|y)$ is designed as a mapping that is independent of $\theta$.
@@ 4.1 Perfect Perception Estimator
Definition 4.1 (Perfect Perception Recovery).
@@
It has been established in previous work [Blau:2018] that a perfect perceptual quality estimator obeys $p_{\hat{X}}(x)=p_{X}(x)$. In [Blau:2018], Blau et al. (2018) also showed that perceptual quality and distortion are at odds with each other, and that for the square error distortion $\|\hat{x}-x\|^{2}$, perfect perceptual quality need not sacrifice more than 3dB in PSNR.
@@ 4.2 Theoretical Analysis
Consider M classes $\theta_{i}$ with the conditional probability density functions (pdf’s) $p(\mathbf{x}|\theta_{i})$ defined on the observation space $\mathcal{X}^{n}$, with priors $P(\theta_{i})$, respectively, i=1,,Mi=1,...,M. The posterior for $\theta_{i}$ is
(18) $q_{i}(\mathbf{x})\triangleq p(\theta_{i}|\mathbf{x})=\frac{P(\theta_{i})p(\mathbf{x}|\theta_{i})}{p(\mathbf{x})},$
where $p(\mathbf{x})$ is the mixture $p(\mathbf{x})=\sum^{M}_{i=1}P(\theta_{i})p(\mathbf{x}|\theta_{i})$. We note that $0\leq q_{i}(\mathbf{x})\leq 1\forall\mathbf{x},\ 1\leq i\leq M$.
@@
(19) $\Gamma_{i}=\Big\{\mathbf{x}\in\mathcal{X}^{n}\ \Big|\ \sum^{M}_{k=1}c_{ik}P(\theta_{k})p(\mathbf{x}|\theta_{k})<\sum^{M}_{k=1}c_{lk}P(\theta_{k})p(\mathbf{x}|\theta_{k}),\ l=1,...,M,l\neq i\Big\},$
@@
(21) $P(\theta_{1})p(\mathbf{x}|\theta_{1})<P(\theta_{2})p(\mathbf{x}|\theta_{2})\quad=>\quad\mathbf{x}\sim p_{\theta_{2}}(\mathbf{x}),$
@@
(22) $R_{2}^{*}=P(\theta_{2})\int_{\Gamma_{1}}p(\mathbf{x}|\theta_{2})d\mathbf{x}+P(\theta_{1})\int_{\Gamma_{2}}p(\mathbf{x}|\theta_{1})d\mathbf{x},$
@@
(23) $\Gamma_{1}=\Big\{\mathbf{x}\in\mathcal{X}^{n}\ \Big|\ P(\theta_{2})p(\mathbf{x}|\theta_{2})<P(\theta_{1})p(\mathbf{x}|\theta_{1})\Big\},$
@@
(24) $\Gamma_{2}=\Big\{\mathbf{x}\in\mathcal{X}^{n}\ \Big|\ P(\theta_{1})p(\mathbf{x}|\theta_{1})<P(\theta_{2})p(\mathbf{x}|\theta_{2})\Big\}.$
Lissack et al., 1976 [Lissack_1976] defined a separability measure between two classes,
(25) $J^{x}_{\alpha}(\theta_{1},\theta_{2})=E\phi_{\alpha}(\mathbf{x}),$
@@
(26) $\phi_{\alpha}(\mathbf{x})=|q_{1}(\mathbf{x})-q_{2}(\mathbf{x})|^{\alpha}\quad 0\leq\alpha\leq\infty.$
@@
(27) $P_{e}=\frac{1}{2}[1-J_{1}(\theta_{1},\theta_{2})].$
@@
$P_{e}$ $=E_{x}\ \mathrm{min}(q_{1}(\mathbf{x}),q_{2}(\mathbf{x}))=\int\mathrm{min}(q_{1}(\mathbf{x}),q_{2}(\mathbf{x}))p(\mathbf{x})d\mathbf{x}$
$=\frac{1}{2}\int\big(1-|q_{1}(\mathbf{x})-q_{2}(\mathbf{x})|\big)p(\mathbf{x})d\mathbf{x}=\frac{1}{2}[1-J_{1}(\theta_{1},\theta_{2})].\@qedbox{}$
The bounds for $\alpha\neq 1$ and for $M>2$ is provided in Lissack et al., 1976 [Lissack_1976]. In other words, the probability of misrecognition is inherently bounded by the absolute value of the difference between the posterior pdfs.
@@
(28) $J^{x}_{1}(\theta_{1},\theta_{2})=\int\big|P(\theta_{2})p(\mathbf{x}|\theta_{2})-P(\theta_{1})p(\mathbf{x}|\theta_{1})\big|d\mathbf{x}.$
@@
(29) $J^{x}_{1}(\theta_{1},\theta_{2})=\int\big|\beta p(\mathbf{x}|\theta_{1})-(1-\beta)p(\mathbf{x}|\theta_{2})\big|d\mathbf{x}.$
@@
(30) $J^{y}_{1}(\theta_{1},\theta_{2})=\int\big|\beta p(\mathbf{y}|\theta_{1})-(1-\beta)p(\mathbf{y}|\theta_{2})\big|d\mathbf{y}.$
@@
(31) $p(\mathbf{y}|\theta)=p_{\theta}(\mathbf{y})=\int p_{\theta}(\mathbf{x},\mathbf{y})d\mathbf{x}=\int p_{\theta}(\mathbf{x})p(\mathbf{y}|\mathbf{x})d\mathbf{x},$
@@
(32) $J^{y}_{1}(\theta_{1},\theta_{2})$ $=\int\Big|\beta\int p(\mathbf{y}|\mathbf{x})p(\mathbf{x}|\theta_{1})d\mathbf{x}-(1-\beta)\int p(\mathbf{y}|\mathbf{x})p(\mathbf{x}|\theta_{2})d\mathbf{x}\Big|d\mathbf{y}$
@@
where the inequality follows from $\int|g(x)|dx\geq|\int g(x)dx|$, and the last equality follows from $\int p(\mathbf{y}|\mathbf{x})d\mathbf{y}=1$. In other words, a classifier operating in the $Y$ domain has larger or equal minimal probability of error as the classifier operating in the $X$ domain. Formally, denote the error of the Bayes classifier operating on the signal $\mathbf{v}$ as $P_{e}(\mathbf{v})$. Thus, according to our derivation we now have,
Theorem 4.2 (Bayes error bound following signal transformation).
@@
(34) $J^{\hat{\mathbf{x}}}_{1}(\theta_{1},\theta_{2})=\int\Big|\int p(\hat{\mathbf{x}}|\mathbf{y})\big[\beta p(\mathbf{y}|\theta_{1})d\mathbf{y}-(1-\beta)p(\mathbf{y}|\theta_{2})\big]d\mathbf{y}\Big|d\hat{\mathbf{x}}\leq J^{y}_{1}(\theta_{1},\theta_{2})\leq J^{x}_{1}(\theta_{1},\theta_{2}).$
@@
Theorem 4.3 (Bayes error bound following class-agnostic signal recovery).
@@
(35) $P_{e}(\hat{\mathbf{x}})\geq P_{e}(\mathbf{y})\geq P_{e}(\mathbf{x}).$
@@
Theorem 4.4 (Bayes error bound following perfect conditional perception recovery).
@@
(37) $J^{\hat{x}}_{1}(\theta_{1},\theta_{2})=\int\Big|\beta\int p(\mathbf{x}|\theta_{1})p(\hat{\mathbf{x}}|\mathbf{x},\theta_{1})d\mathbf{x}-(1-\beta)\int p(\mathbf{x}|\theta_{2})p(\hat{\mathbf{x}}|\mathbf{x},\theta_{2})d\mathbf{x}\Big|d\hat{\mathbf{x}}$.
@@
(38) $\arg\max_{\mathbf{x}}p_{\theta}(\mathbf{x},\mathbf{y})=\arg\max_{\mathbf{x}}p(\mathbf{x}|\mathbf{y})p(\theta|\mathbf{x})$
@@
Related Work. Jalal et al., 2021 [Jalal:2021] state that in light of the “White Obama” controversy (Menon et al., 2020b), it has been suggested that reconstruction algorithms are biased because the datasets used for training and inference are unbalanced, thus the obtained models misrepresent the true population distribution. They [Jalal:2021] also claim that reconstruction algorithms that are designed to maximize accuracy are inclined to increase bias. They define proportional representation (PR) such that given a collection of sets of images $\{\mathcal{X}_{\theta_{i}}\}_{i=1}^{K}$ where $\theta_{i}$ is a vector of sensitive attributes and each $\mathcal{X}_{\theta_{i}}$ represents the group carrying the sensitive attributes $\theta_{i}$,
(39) $P(\hat{\mathbf{x}}\in\mathcal{X}_{\theta_{i}})=P(\mathbf{x}\in\mathcal{X}_{\theta_{i}})\ \forall i.$
@@
(40) $P(\hat{\mathbf{x}}\in\mathcal{X}_{\theta_{i}}|\mathbf{y})=P(\mathbf{x}\in\mathcal{X}_{\theta_{i}}|\mathbf{y})\ \forall i.$
They show that randomly sampling from the posterior $\hat{\mathbf{x}}\sim p(\mathbf{x}|\mathbf{y})$ achieves CPR and PR. Hence, fairness is accomplished because every image that could have led to the measurement could be represented, and not just the most likely one. Note that by these definitions the reconstruction is stochastic, and not deterministic, and that any reconstruction that falls outside the set $\mathcal{X}_{\theta_{i}}$ is considered a wrong outcome. Furthermore, these definitions ignore the distribution $p_{\theta_{i}}(\mathbf{x})$. Thus, Ohayon et al. [Ohayon:2024b] propose other measures of fairness to improve the estimation of the distribution of the different groups. However, note that, in our case study, our objective is to estimate $\theta_{i}$ to the best accuracy. We are not seeking to improve fairness nor to provide a more diverse distribution of the reconstruction $\mathbf{x}$.
@@ 4.3 Examples
In this subsection we provide illustrative examples to enhance the readers intuitive understanding of the discrete problem setting, and of our proposed theoretical analysis.
@@ 4.3.1 Naive tree example
Consider the setting partially depicted in figure 7 such that, $\theta=\{\theta_{1},\theta_{2}\}$ represent two classes with probabilities $P(\theta_{1})=0.8,\ P(\theta_{2})=0.2$. The $x$ values in each class are $A_{X}=\{0,1,2\}$, $B_{X}=\{3,4,5\}$, which are conditionally equiprobable, $\big\{x\in A_{X}:p_{\theta_{1}}(x)=\big[\frac{1}{3},\frac{1}{3},\frac{1}{3}\big]\big\}$, $\big\{x\in B_{X}:p_{\theta_{2}}(x)=\big[\frac{1}{3},\frac{1}{3},\frac{1}{3}\big]\big\}$. The $y$ values in each class are $A_{Y}=\{0.5,1.5,2.5\}$, $B_{Y}=\{1.5,3.5,4.5\}$, and the conditional probabilities of observing $y=1.5$ are p(y=1.5|x=4)=1,p(y=1.5|x=1)=0.4p(y=1.5|x=4)=1,\ p(y=1.5|x=1)=0.4. Namely, the original sets are disjoint but the observation sets partially coincide, and one set is 4 times more likely than the other.